Find hcf of 65 and 117 and express in the form 65m+ 117n.

Grade:10

1 Answers

Utkarsh Sharma
6 years ago
65=5×13117=3×3×13So we see 13 is the common factor 13 is the HCFNow let n be equal to -1 and m equal to 2So 65m+117n65(2) + 117(-1)=13(HCF)130-117=13(HCF)